OpenStudy (adll23):

FAN AND MEDAL 1. Solve the following system of equations: (1 point) -2x + y = 1 -4x + y = -1 (3, 1) (-1, 3) (-1, -3) (1, 3)

OpenStudy (photon336):

well the idea is that you need to eliminate one of the variables by combining the two equations together.

OpenStudy (photon336):

here's what you do: you see that in both equations you have +y right?

OpenStudy (adll23):

yes

OpenStudy (adll23):

would you take away one y?

OpenStudy (photon336):

let's chose to get rid of y. what number if multiplied by y would result in both y's canceling out or resulting in their sum being zero

OpenStudy (adll23):

1?

OpenStudy (photon336):

almost it's -1 you see why?

OpenStudy (adll23):

oh cause it is positive!

OpenStudy (photon336):

okay so now. check this out

OpenStudy (photon336):

we must multiply one of the equations by -1. everything because we need to ensure that when we do something to an equation it's done to everything let's pick the top equation we multiply everything by -1 -1*(-2x+y = 1) becomes = -1*-2y+ (-1*y) = -1*1 2x-y = -1

OpenStudy (adll23):

oh i see now do we do that to the other equation too?

OpenStudy (photon336):

now we get this: 2x-y = -1 -4x+y = -1 Now the point of this is that when we ADD the two equations together y cancels out. so try adding these equations together and then tell me what you get?

OpenStudy (adll23):

-2x=-2?

OpenStudy (photon336):

yep :)

OpenStudy (photon336):

so now, you know what x is

OpenStudy (adll23):

and then do you divide by -2?

OpenStudy (photon336):

yep so what's our x value?

OpenStudy (adll23):

-1?

OpenStudy (photon336):

remember -2x = -2 x = 1 because -2*1 = -2

OpenStudy (adll23):

ohhh yeah

OpenStudy (adll23):

so D is the correct answer?

OpenStudy (photon336):

now how do you think we would find Y?

OpenStudy (photon336):

We're almost done

OpenStudy (adll23):

by putting 1 into the equation?

OpenStudy (photon336):

so let me show you something: we know that x = 1 right? to find y all we need to do is plug in x =1 into one of the equations -2x+y = 1 try doing that and showing me what you get.

OpenStudy (adll23):

-2(1)+y=1 -2+y=1 +2 +2 y=3 is this right?

alones (alones):

so you got 3 for y?

OpenStudy (photon336):

=) yep

OpenStudy (photon336):

@adll23 the most important step is eliminating the variable

OpenStudy (adll23):

yes it is:) YAYAYAY thank you so much! you made it so much easier for me!

OpenStudy (photon336):

first ask yourself what number must i multiply one equation by such that when I add them both together, one variable cancels out

OpenStudy (adll23):

kk

OpenStudy (photon336):

Also multiplying a number by an entire equation is kind of like the same thing you aren't doing something to it. so if you multiply one number say y by -1 it has to be done to the entire equation that's all. :)

OpenStudy (photon336):

well i should mean you aren't changing anything by doing that

OpenStudy (adll23):

yes! thank you so much for the help, Photon:)

OpenStudy (photon336):

no problem anytime :)

OpenStudy (photon336):

if you want you can try to find (x and y) in this problem 2x+y = 3 x+y = -2